OpenStudy (anonymous):

solve please t/6 + t/8 = 1

OpenStudy (dumbcow):

add fractions, find common denominator 8*3 = 24 6*4 = 24

OpenStudy (anonymous):

4t/24 + 3t/24 = 7t/24 = 1 7t=24 t=24/7